There and some benches in a classroom. If 4 students sit on it bench then 3 benches are left on occupied and each 3 students sit on each bench, then 3 students are left standing. How many students are there in the classroom?

Grade:11

1 Answers

Arun
5 years ago
Dear Bipin

Let there be x benches.
Then , number of students = 4(x-3) = 3(x) + 3
x = 15
students = 4(x-3) = 48
 
Regards
Arun (askIITians forum expert)